Fermi National Laboratory


Accelerator Update

Friday March 5
The day shift began with Operations monitoring store 3273, with Antiproton Source (Pbar) stacking, and with Main Injector (MI), Recycler, and MiniBooNE experts conducting studies.

At 9:38 AM, Operations turned off the MI beam while a RF technician worked on station MIRF13. He completed his work and stacking resumed at 10:07 AM.

Control techs began replacing a CAMAC power supply for Pbar at 10:35 AM. They completed their work in a little less than an hour.

At 11:12 AM, the Tevatron (TeV) suffered a sector F4 quench. TeV experts had just begun a study period. The quench was due to their studies. The D-zero (D0) experimenters accessed their collision hall.

A Linac RF station (LRF2) tripped off at 5:23 PM and wouldn't reset. Operators called in a Linac tech and he managed to get the station running normally at 6:44 PM. Stacking resumed.

D0 came out of access at 6:49 PM.

Operation put the TeV into shot setup at 8:52 PM. At 11 PM, Operations established store 3275 with an initial luminosity of 65.77E30.

Saturday March 6
The midnight shift began with Operations monitoring store 3275, with Pbar stacking, and with MiniBooNE taking beam.

At 4:29 AM, Recycler experts began stashing antiprotons from Pbar. By 5:50 AM, Recycler had stashed 48mA.

Pbar resumed stacking at 5:57 AM.

At 8:12 AM, operators noticed that the beam intensity on the Pbar lithium lens target had dropped. Operators began tuning Booster.

Operations established beam to the Switchyard (SY) dump at 9:56 AM and then began working on sending beam to Meson Center (MCenter). Two hours later, with the help of an expert, Operations had beam to MCenter.

Do reported at 10:10 AM that the electrical noise that had been troubling them was gone. They had raised the cooling water temperature, lowered the hall humidity, and reconnected part of the toroid, but that doesn't mean the problem won't come back.

LRF2 tripped off at 10:35 AM. It reset okay.

While a Booster expert was conducting studies, operators noticed that a Pbar kicker (A:IKIK) stopped pulsing at 2:35 PM and wouldn't come back on. Operators called in an expert. Another expert saw that the ramp had been disabled. Everything was reset. Investigation continued, but no clear cause was found.

Booster studies ended at 3:33 PM.

A MI/NuMI study began at 5:35 PM.

At 5:42 PM, due to an alarm indication of low oxygen the Collider Detector at Fermilab (CDF) had its detector's high voltage trip off. A cryo system tech investigated and found it to be a false indication. CDF experts managed to get their detector powered up by 7:36 PM.

A Linac RF station (LRF4) tripped off and then tripped off twice more while operators tried to run it up. At 7:37 PM, operators got the station back on.

At 10:13 PM, Operations turned off the Booster notcher for about a half hour to let NuMI experts complete their studies. Booster was back to normal by 10:48 PM.

Sunday March 7
The midnight shift began with Operations monitoring store 3275, with Pbar stacking, and with beam to MiniBooNE.

At 12:11 AM, Operations turned off beam to MCenter. The experimenters accessed their experimental hall.

Two Linac RF station (KRF3 & 7) tripped off at 2:52 AM. Upon further investigation, operators discovered that the Tevatron Electron Lens (TEL), among other devices, also tripped off at the same time. Operators thought the cause may have been due to a small power glitch. Operations had most device reset by 3 AM.

LRF2 tripped off at 4:03 AM. After numerous attempts, operators got the station up and running. But for the next hour and a half this station continued to trip, sometimes causing LRF1 to trip off also. At 5:37 AM, LRF2 tripped off and would not reset. Operators called in a Linac tech. He got the station back on by 6:10 AM, but reported that there were more problems. He took the station down at 6:47 AM, made repairs and had it back in service twenty minutes later. Stacking resumed.

Operations began working on reestablishing beam to MCenter at 9:21 AM. Within an hour they had beam to the experiment.

A TeV TEL expert began an end-of-store study at 9:52 AM.

At 10:27 AM, a Booster expert began a cogging study.

Operations terminated store 3275 at 10:34 AM.

CDF experimenters accessed their collision hall at 10:52 AM. They came out of access an hour and a half later.

At 1:10 PM, Operations put the TeV into shot setup.

At 2:47 PM, a Booster power supply (ORBMP) tripped off. Operators investigated and could smell burned electronics. They called in two Booster experts. The experts discovered a blown SCR bank and related equipment. Another expert came in to help with the repairs. The following is a picture of a blown snubber board.

Operations established store 3277 at 2:52 PM with an initial luminosity of 58.82E30.

Pbar resumed stacking at 5:55 PM.

Operations reestablished beam to MCenter at 5:58 PM.

At 6:07 PM, LRF2 tripped off for the first of three times in five minutes.

At 6:17 PM, a Pbar Debuncher dipole power supply (D:IB) tripped off. An expert came in at 6:54 PM and replaced the phase imbalance circuit on QT405.

Booster suffered a RF section vacuum burst at 6:36 PM. Everything reset okay.

Operations turned off beam to MCenter and SY at 7:26 PM. MCenter experimenters went in on access.

At 8:04 PM, Recycler lost its stash of antiprotons. Experts investigated and found a correction magnet power supply had tripped off. Techs had this supply repaired by 9:07 PM.

Recycler experts began stashing antiprotons at 10:58 PM.

Stacking resumed at 11:43 PM.

Monday March 8
The midnight shift began with Operations monitoring store 3277, with Pbar stacking, and with MiniBooNE taking beam.

At 2:39 AM, CDF requested Operations to tune the TeV beam.

Today's Plans
The plans for today start with allowing MI, Recycler, SY, and MiniBooNE to conduct studies, to continue monitoring store 3277, and later this morning to allow techs to access SY to work on a septum (FSEP).

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 3/8/2004   email Fermilab